Sol Summary Report – Mar 25th

Summary Title : Day #9 (Out of Sim)

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

COM continued data collection for the thoracentesis and bike-powered 3D printing studies.

COM + XO continued 3D printing items in Level 1 of the Hab.

ENG + GHO signed out their duties to XO + COM.

ENG + GHO conducted testing of the new drone controller, completed surveys, and submitted qualitative feedback to manufacturer.

XO transported ENG/Astronomer + GHO/Journalist to GJT this PM.

Look ahead plan : Continue 3D printing projects, drone flight mapping, drone controller testing, and data collection for bike-powered 3D printing studies.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : Drone flight testing in PM.

Reports to be filed : Operations, GreenHab, Astronomy

Support Requested : None

Sol Summary Report – Mar 24th

Summary Title : Day #8 (Out of Sim)

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

Last night, Crew 207 successfully completed our check-out inspection with David. COM continued data collection for the thoracentesis study.

This AM XO transported HSO + SCI to GJT. ENG + GHO hiked up to the North Ridge to observe fossils.

COM continued 3D printing items in Level 1 of the Hab.

Look ahead plan : Continue drone flight mapping + data collection for thoracentesis and bike-powered 3D printing studies. COM to begin inputting thoracentesis study data values for remote analysis and manuscript preparation. XO to transport ENG + GHO to GJT in PM on Sol 9.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : Drone mapping flights in AM out of sim.

Reports to be filed : Operations, GreenHab

Support Requested : None

Sol Summary – March 23rd

Summary Title: Day #7 of Sim

Author: Julielynn Wong

Mission Status: All nominal

Sol Activity Summary:

Yesterday evening, COM continued data collection for the thoracentesis study.

This morning’s EVA successfully completed two additional drone mapping flights of MDRS. This afternoon’s EVA completed flight testing of a square (60 feet) flight pattern at an altitude of 20 feet using a traditional and new drone controller by two drone pilots. The third pilot conducted a habitat roof inspection with the new controller.

A post-flight survey was completed.

HSO + SCI continued 3D printing items in Level 1 of the Hab and tested their 3D printed labware in the Science Dome.

Look ahead plan: Clean up MDRS this evening for inspection. We will break sim tomorrow and transport two crew members (HSO + SCI) to GJT on Sol 8. ENG + GHO to depart on Sol 10. Mission summary to follow later on this week.

Anomalies in work: None

Weather: Sunny

Crew physical status: Fine

EVA: AM: Drone mapping PM: Drone controller testing

Reports to be filed: Operations, GreenHab, Journalist, EVA Report, Research Report

Support Requested: None

Sol Summary- March 18th

Summary Title : Day #1 of Sim

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary: ENG prepared breakfast. COM and HSO set up a research study in the Science Dome. HSO and SCI planned tomorrow’s EVA. GHO prepared lunch using cilantro and basil from the Greenhab. Weather was clear and sunny for EVA #1 to Zubrin’s Head and Kissing Camels. AST continued operations in the solar observatory. GHO made dinner.

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Environmental sample container

Toothbrush stand/holder

Selfie stick with attached rearview mirror for EVAs

Conductive stylus for EVAs

Headband with scratcher attachment for EVAs

Contour map for road to MDRS

Look ahead plan : We are planning another EVA to scout out other locations suitable for drone mapping. We are preparing files to be 3D printed at MDRS to save time and money for future crews.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA : Zubrin’s Head + Kissing Camels

Reports to be filed : Operations, GreenHab, Journalist, Mission Plan, EVA Report, EVA Request

Support Requested : David generously loaned Crew 207 his manual can opener.

Sol Summary – March 22nd

Summary Title : Day #6 of Sim

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

Yesterday evening, COM continued data collection for the thoracentesis study. Last night, Crew 207 discharged the battery charged by a bike by running 3D4MD’s 3D printing system for an additional 30 minutes.

This morning’s EVA successfully completed two out of four drone mapping flights of MDRS. As well, one Part 107 drone pilot successfully piloted a camera drone to conduct an outdoor inspection of the Hab roof.

This afternoon’s EVA completed flight testing of a square (60 feet) flight pattern at an altitude of 20 feet using a traditional and new drone controller by three drone pilots. A post-flight survey was completed.

Crew 207 continued 3D printing items in Level 1 of the Hab. We all continued to use our 3D printed straws to identify and reuse our drinking cups and thereby, converse water consumed by washing dishes. We also 3D printed tongue depressors (which are included in the ISS medical inventory) out of food-safe, biodegradable plastic using a myriad of colours. We had a brief power interruption while 3D printing a slide drying rack out of washable, biodegradable red plastic but luckily, the 3D4MD 3D printing system was able to continue and finish the print after the power was restored.

Crew 207 identified the following useful item that could be 3D printed out at MDRS:

Earpiece/microphone wire clip

We took a photograph of a Senior Member to send a greetings from “Mars” to the Civil Air Patrol cadets of Niagara Falls #1 Composite Squadron. In the photo, a Senior Member is holding a 3D printed coat hanger which was printed on demand locally out of biodegradable plastic at MDRS.

Crew 207 members are exercising on our bike to re-charge a battery to power 3D4MD’s 3D printing system.

Look ahead plan : We will conduct two walking EVAs tomorrow to continue our drone mapping and controller testing. We will continue to 3D print items at MDRS to save time and money for future crews. We will continue data collection for our thoracentesis and bike-powered 3D printing projects.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : AM: Drone mapping PM: Drone controller testing

Reports to be filed : Operations, GreenHab, Journalist, EVA Reports, EVA Request

Support Requested : None

Sol Summary – March 20th

Summary Title : Day #4 of Sim

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary: GHO test rode the bike-powered 3D printing apparatus for 1 hour in Level 1 of the Hab last night. Last night and today, COM collected study data for the thoracentesis project in the Science Dome. Crew 207 continued 3D printing systems in Level 1 of the Hab. AST continued operations in the solar observatory. Weather was clear and sunny with gusting winds for EVA #3. EVA Team completed a post-EVA human factors survey on using the Mavic Pro standard drone controller.

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Headset ear piece

Look ahead plan : We will conduct a walking EVA tomorrow to continue our drone flight testing. We will continue to 3D print items at MDRS to save time and money for future crews. We will continue data collection for our thoracentesis and bike-powered 3D printing projects.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA :

Reports to be filed : Operations, GreenHab, Journalist, EVA Report, EVA Request

Support Requested : David to provide Crew 207 with a map showing the property boundaries of MDRS

Sol Summary – March 19th

Summary Title : Day #2 of Sim

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary: XO arrived at Hab with the remainder of our research equipment. Crew 207 set up 3D printing systems in Level 1 of the Hab and RAM and set up the bike-powered 3D printing apparatus in Level 1 of the Hab. COM prepared the second part of thoracentesis project in the Science Dome. ENG checked the status of EVA suits. XO provided a fresh battery for the RAM smoke detector. Weather was clear and sunny for EVA #2 to Cattle Gate and Zubrin’s Head. XO and GH planned tomorrow’s EVA. AST continued operations in the solar observatory.

SCI and HSO 3D printed the following labware items out of
biodegradable, food-safe plastic:

Microscope cover slide

Western blot comb

Soil sieve

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Shade cover of mobile devices (including iPad, iPhone + GPS)

Look ahead plan : We will conduct a walking EVA to scout out a location suitable for drone mapping of MDRS. We will continue to 3D print items at MDRS to save time and money for future crews. We will initiate data collection for our thoracentesis and bike-powered 3D printing projects.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA : Cattle Gate, Zubrin’s Head

Reports to be filed : Operations, GreenHab, Journalist, EVA Report, EVA Request, Astronomer

Support Requested : None

Sol Summary – March 17th

Summary Title: Starting our Sim

Author: Julielynn Wong

Mission Status: All nominal

Sol Activity Summary: Crew 206 departed this morning. Crew 207 Greenhab Officer prepared breakfast. Crew 207 Commander submitted yesterday’s food inventory. Crew 207 HSO conducted a safety check. Crew 207 Scientist drafted the cooking and cleaning schedule. Commander cleaned and organized both levels of the Hab and the Science Dome. Greenhab Officer and Crew 207 Engineer/Astronomer began planning tomorrow’s EVA. Crew Scientist prepared lunch with fresh basil seasoning from the Greenhab. Engineer/Astronomer began operations in the solar observatory. Crew 207 completed paperwork and training with the MDRS intern. All 5 crew members successfully test drove the rovers under the supervision of the MDRS intern. Crew 207 completed their EVA spacesuit training.

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Coat hanger

Bag clip

Soap holder

S hook

Stackable storage bin

Pen storage container

Funnels

Labware drying rack

Soil sieves

Water dipstick

Selfie stick

Dishes

Colored straws to help crew members identify their individual cups

Scratch stick for the inside of the EVA spacesuit helmets

Look ahead plan: We are planning an EVA to scout out a location for drone flight testing. We are preparing files to be 3D printed at MDRS to save time and money.

Anomalies in work: None

Weather: Sunny all day

Crew physical status: Fine

EVA: N/A

Reports to be filed: Operations, GreenHab, HSO, Scientist, Journalist, EVA Request

Support Requested: None

Sol Summary – March 16th

Crew 206 March 16th, 2019
Sol 20

Summary Title: Welcoming Crew 207

Author: Jeremy Auclair

Mission Status: All nominal

Sol Activity Summary: Yesterday night we invited Atila and David to have some homemade pizza at the Hab, it was really nice to see some new faces. This morning we woke up earlier than usual, around 06:00 am, to watch the sunrise from Hab Ridge. The view is breathtaking. It felt great to feel the cold breeze on our skins. We stayed up there for more than 2,5 hours. After that, we ate breakfast and got ready for our hike.

At 10:00 am we took Spirit, Opportunity, and Curiosity to go towards the Copernicus road. It was a beautiful hike, we walked and sunbathed for a good amount of time. We didn’t have any time constraints because the last news we had was that Crew 207 would only arrive tonight around 08:00 pm.

We came back at the Hab around 01:30 pm, and at this point, Atila gave us the news that Crew 207 would arrive in an hour. We had to rush to finish cleaning the Hab and prepare lunch. But the crew arrived later and we ate alone.

The handover went really well, we trained them with the time we had and talked a little about our research and our lives. We will go to Stan’s Burger around 08:00 pm so the COMMS window has to be managed beforehand.

Look ahead plan: Tomorrow we will get up at 06:00 am to drive back to Grand Junction, rent a car and head straight to Denver, it’s gonna be a long day.

Anomalies in work: None

Weather: Sunny all day

Crew physical status: Fine

EVA: Copernicus exploration

Reports to be filed: GreenHab, Journalist, Operation, EVA Request, EVA Report

Support Requested: None

Sol Summary Report – Mar 15th

Crew 206 March 15th 2019
Sol 19

Summary Title : Wrapping up the Sim

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Today is our last “routine” day. We gave everything we had for the workout session, we ate quickly to be ready on time for the EVA which was earlier than usual. Half of the crew went exploring the ends of Galileo road while the other half started packing and continued cleaning the Hab, with only a few rest moments in the morning. The EVA teamed wrapped up the Weather station and LOAC disassembly around 10:30 am, and after they had arranged their equipment we started cleaning the lower deck. Everyone has a lot to do today, and we kept this fast pace all day long. At 11:30 am we started working on the daily reports and the preparation of lunch. After a copious meal, we went all around the station to pack our small equipment, film some last scenes and gather the last experimental
data. Even on this busy day we had to do the TELEOP session and the VR experiment, for the last time. The important end of mission reports are already done, so we are a little ahead of schedule. After the VR session, at 18h00, we opened the main airlock to breathe some fresh and absorb the sunlight. This is the end, we stayed outside for more than 45 minutes, taking in the surroundings for one last time. We are happy of the work we did, the experiences we lived and the memories we formed.

Look ahead plan : We are looking forward to meeting Crew 207, it will feel good to meet some new faces after 3 weeks of isolation. There is still work to do before leaving but the vast majority has been done. We will keep some personal time tomorrow morning to watch the sunrise from the Hab ridge, and perhaps jog towards the Mesa.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA : Weather station and LOAC disassembly and Galileo road exploration

Reports to be filed : GreenHab, Journalist, Operation, EVA Request, EVA Report

Support Requested : None

Copyright © The Mars Society. All rights reserved. | Main Site